OpenStudy (anonymous):

Can you check my work? v(t)=t^3-2t^2+1 s(0)=1 Find the position function of the particle

OpenStudy (anonymous):

\[\int\limits_{0}^{t}(t^3-2t^2+1)dt=s(t)-0\]

OpenStudy (anonymous):

s(t)-1

OpenStudy (anonymous):

\[\frac{ t^4 }{ 4 }-\frac{2t^3 }{ 3 }+t-1=s(t)-1\]

OpenStudy (anonymous):

\[s(t)=\frac{ t^4 }{ 4 }-\frac{ 2t^3 }{ 3 }+t\]

OpenStudy (anonymous):

that is right

OpenStudy (anonymous):

but if i use indefinite integral i get s(t)=t^2/2-2t^3/3+t+1

OpenStudy (anonymous):

v(t) is the velocity if you integrate that you would get the position ( you do not need limits from 0 to t)

OpenStudy (anonymous):